OpenStudy (anonymous):

PLZ HELP graphing hyperbola i am only confused on the asymptoes part! (y-7)^2/-(x)^2/33=1

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

\[\frac{ (y-7 )^{2}}{ 4 }-\frac{ x ^{2} }{ 33 }=1\]

OpenStudy (anonymous):

@Hero plz help only the asymptoes part.

OpenStudy (anonymous):

I tried this formula but that is not right http://www.mathwarehouse.com/hyperbola/graph-equation-of-a-hyperbola.php

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

have a look at this image http://www.mathwarehouse.com/hyperbola/images/picture-of-hyperbola-vertical-transverse.gif

jimthompson5910 (jim_thompson5910):

the only real difference between this and what you have is that you would need to shift everything up 7 units

OpenStudy (anonymous):

so whatever my asymptoe is i have to shift it up 7?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes because they're describing asymptotes for the general equation but you've shifted that general equation up 7 units

OpenStudy (anonymous):

@jim_thompson5910 so it is the center??

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

k so \[\frac{ (x-5)^{2} }{ 49 }-\frac{ (y-1)^{2} }{ 17 }=1\]

OpenStudy (anonymous):

so here the asymptope will be at 5,1

jimthompson5910 (jim_thompson5910):

yes they cross at (5,1)

jimthompson5910 (jim_thompson5910):

where the center is, if that's what you're asking

OpenStudy (anonymous):

thank you so sos sos much u have cleared a lot of stuff that was on my mind i don't knwo how to thank you!but thank you!

jimthompson5910 (jim_thompson5910):

you're welcome